The length of a rectangle is 6 meters more than twice the with. If the area of the rectangle is 140 square meters, find the dimensions.
Answer by checkley77(12844)

L=2W+6
A=LW
140=(2W+6)W
140=2W^2+6W
2W^2+6W-140=0
2(W^2+3W-70)=0
2(W+10)(W-7)=0
W-7=0
W=7 ANS.
L=2*7+6
L=14+6
L=20 ANS.
PROOF:
140=7*20
140=140